# A Micro-Macro Decomposition-Based Asymptotic-Preserving Random Feature Method for Multiscale Radiative Transfer Equations

### v1

@@ 2.2 Random Feature Method
which is analogous to the finite element method and one can generalize the partition to a more general shape if necessary. Assume $\bm{y}=(y_{1},\cdots,y_{d})^{T}\in\Omega$, we apply the following transformation to obtain a partition dependent normalized vector $\tilde{\bm{y}}_{i}=(\tilde{y}_{i1},\cdots,\tilde{y}_{id})^{T}$:
@@
where $\bm{w}_{ij}\in\mathbb{R}^{d}$ and $b_{ij}\in\mathbb{R}$ are randomly generated with the uniform distribution in $[-B,B]$ and fixed parameters, and $\sigma$ is a scalar function, named activation function.
@@ 2.3 Difficulty of RFM to resolve small scales
$\mathcal{T}^{\varepsilon}f_{M}=\mathcal{T}^{\varepsilon}\left(\sum_{i=1}^{M}\psi_{i}(x,v)\sum_{j=1}^{J_{i}}f_{ij}\phi_{ij}(x,v)\right)=\sum_{i=1}^{M}\sum_{j=1}^{J_{i}}f_{ij}\mathcal{T}^{\varepsilon}(\psi_{i}\phi_{ij}).$ (23)
@@
$\bm{A}=\begin{bmatrix}\bm{A}_{\text{int}}\\ \bm{A}_{\text{bdy}}\end{bmatrix}\in\mathbb{R}^{N\times Q},\;\bm{b}=\begin{bmatrix}\bm{b}_{\text{int}}\\ \bm{b}_{\text{bdy}}\end{bmatrix}\in\mathbb{R}^{N},\;N=N_{\text{int}}+N_{\text{bdy}},$ (24)
@@
$\bm{A}_{\text{int}}=\begin{bmatrix}s^{\text{int},1}_{11}&\cdots&s^{\text{int},1}_{1J_{1}}&s^{1}_{21}&\cdots&s^{\text{int},1}_{MJ_{M}}\\ s^{\text{int},2}_{11}&\cdots&s^{\text{int},2}_{1J_{1}}&s^{2}_{21}&\cdots&s^{\text{int},2}_{MJ_{M}}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ s^{\text{int},N_{\text{int}}}_{11}&\cdots&s^{\text{int},N_{\text{int}}}_{1J_{1}}&s^{\text{int},N_{\text{int}}}_{21}&\cdots&s^{\text{int},N_{\text{int}}}_{MJ_{M}}\end{bmatrix},\,\bm{b}_{\text{int}}=-\varepsilon\begin{bmatrix}v_{\text{int}}^{1}\\ v_{\text{int}}^{2}\\ \vdots\\ v_{\text{int}}^{N_{\text{int}}}\end{bmatrix},$ (25)
$\bm{A}_{\text{bdy}}=\begin{bmatrix}s^{\text{bdy},1}_{11}&\cdots&s^{\text{bdy},1}_{1J_{1}}&s^{\text{bdy},1}_{21}&\cdots&s^{\text{bdy},1}_{MJ_{M}}\\ s^{\text{bdy},2}_{11}&\cdots&s^{\text{bdy},2}_{1J_{1}}&s^{\text{bdy},2}_{21}&\cdots&s^{\text{bdy},2}_{MJ_{M}}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ s^{\text{bdy},N_{\text{bdy}}}_{11}&\cdots&s^{\text{bdy},N_{\text{bdy}}}_{1J_{1}}&s^{\text{bdy},N_{\text{bdy}}}_{21}&\cdots&s^{\text{bdy},N_{\text{bdy}}}_{MJ_{M}}\end{bmatrix},\;\bm{b}_{\text{bdy}}=\begin{bmatrix}f_{\text{bdy}}^{1}\\ f_{\text{bdy}}^{2}\\ \vdots\\ f_{\text{bdy}}^{N_{\text{bdy}}}\end{bmatrix}.$ (26)
@@
$\bm{f}=\begin{bmatrix}f_{11}&\cdots&f_{1J_{1}}&f_{21}&\cdots&f_{MJ_{M}}\end{bmatrix}^{T},$ (27)
@@
First, we consider the problem (17) with $\varepsilon=1$ to validate the effectiveness of the RFM. Here, we set the number of partitions $M=1$ with the PoU function $\varphi_{a}$. We choose the $\tanh$ function as our activation function and the number of random feature functions $J_{i}=2^{n},n=4,5,6,7$. The weights $\bm{w}_{ij}$ and biases $b_{ij}$ are randomly generated with $B=1$. Besides, the collocation points are uniform grids with $(N_{x},N_{v})=(32,64)$. The integration of the operator $\mathcal{L}$ is approximated by Gauss-Legendre quadrature with $16$ points.
The relative $\ell^{2}$ error of the solution $f_{M}$ concerning the degrees of freedom $Q$ is shown in Figure 2. One can observe that the error decreases exponentially to the degrees of freedom $Q$. We also plot the solution $f_{M}$ obtained by the random feature method with $J_{i}=2^{7}$ and reference solution $f_{\text{ex}}$ in Figure 3. The relative $\ell^{2}$ error of the RFM is $3.99\times 10^{-7}$.
Figure 2: Relative $\ell^{2}$ error of the RFM solution $f_{M}$ with respect to the degrees of freedom $Q$.
@@
Next, we test a sequential of problems (17) with $\varepsilon=10^{-2^{i}}$, $i=1,2,3,4$ with the same setting as the previous problem except for the number of random features $J_{i}$ and the collocation points $(N_{x},N_{v})$. We record the relative $\ell^{2}$ error of the RFM solution to the degrees of freedom and the number of collocation points in Table 1 and Table 2, respectively. One can find that the small-scale problems ($\varepsilon\ll 1$) are much more difficult to resolve than the large-scale problems in terms of the degrees of freedom and the number of collocation points.
Table 1: The dependency of the RFM on degrees of freedom. The collocation points are uniform grids with $(N_{x},N_{v})=(128,256)$.
$\varepsilon$ $J_{i}$
@@
Table 2: The dependency of the RFM on the number of collocation points. The number of random feature functions $J_{i}=128$.
@@
To resolve the small-scale information $\mathcal{O}(\varepsilon^{2})$, an increase in both the degrees of freedom and the number of collocation points is required. This leads to a large matrix $\bm{A}$ with a high condition number. The high condition number of the matrix $\bm{A}$ will make the least squares problem (28) ill-conditioned, and the solution $\bm{f}^{*}$ may not be accurate enough. The above analysis implies that the vanilla RFM based on least squares formulation for the radiative transfer equation (1) has tremendous difficulty in resolving small scales in the radiative transfer equation. This is not only due to the ill-conditioning of the least squares problem, but also the lack of capturing the small-scale part of the solution $f(x,v)$. Such a similar phenomenon is also observed in the numerical experiments in [28, 27] when solving the multiscale kinetic equations with small parameters through deep neural networks.
@@ 3 Asymptotic-Preserving Random Feature Method
Assume $\mathcal{F^{\varepsilon}}$ is the microscopic equation that depends on the scale parameter $\varepsilon$ and $\mathcal{F}^{0}$ is its macroscopic limit equation as $\varepsilon\to 0$. The latent solution of $\mathcal{F^{\varepsilon}}$ is approximated by the RFM and solved by the least squares minimization problem $\mathcal{R}(\mathcal{F^{\varepsilon}})$. The asymptotic limit of $\mathcal{R}(\mathcal{F^{\varepsilon}})$ as $\varepsilon\to 0$, if exists, is denoted by $\mathcal{R}(\mathcal{F}^{0})$. If $\mathcal{R}(\mathcal{F}^{0})$ is a “good” least squares formulation of $\mathcal{F}^{0}$, which can capture the multiscale information, then the method is called asymptotic-preserving.
@@
Therefore, equation (34) and (36) constitute the micro-macro system of the radiative transfer equation (31):
@@
respectively. Here, $\{(\bm{x}_{\text{int}}^{k},\bm{v}_{\text{int}}^{k})\}_{1\leq k\leq N_{\text{int}}}$ and $\{(\bm{x}_{\text{bdy}}^{k},\bm{v}_{\text{bdy}}^{k})\}_{1\leq k\leq N_{\text{bdy}}}$ are the interior and boundary collocation points, and $\{f_{\text{bdy}}^{k}\}_{1\leq k\leq N_{\text{bdy}}}$ are the boundary values. $\{\lambda_{\text{int},1}^{k}\},\{\lambda_{\text{int},2}^{k}\}$ and $\{\lambda_{\text{bdy}}^{k}\}$ are regularization parameters. The degrees of freedom of the APRFM are $Q=\sum_{i=1}^{M^{\rho}}J_{i}^{\rho}+\sum_{i=1}^{M^{g}}J_{i}^{g}$.
@@
$\mathcal{T}_{11}^{\varepsilon}\rho_{M}=\sum_{i=1}^{M^{\rho}}\sum_{j=1}^{J_{i}^{\rho}}\rho_{ij}\mathcal{T}_{11}^{\varepsilon}(\psi_{i}^{\rho}\phi_{ij}^{\rho}),\;\mathcal{T}_{12}^{\varepsilon}g_{M}=\sum_{i=1}^{M^{g}}\sum_{j=1}^{J_{i}^{g}}g_{ij}\mathcal{T}_{12}^{\varepsilon}(\psi_{i}^{g}\phi_{ij}^{g}),$ (47)
@@
$\mathcal{T}_{21}^{\varepsilon}\rho_{M}=\sum_{i=1}^{M^{\rho}}\sum_{j=1}^{J_{i}^{\rho}}\rho_{ij}\mathcal{T}_{21}^{\varepsilon}(\psi_{i}^{\rho}\phi_{ij}^{\rho}),\;\mathcal{T}_{22}^{\varepsilon}g_{M}=\sum_{i=1}^{M^{g}}\sum_{j=1}^{J_{i}^{g}}g_{ij}\mathcal{T}_{22}^{\varepsilon}(\psi_{i}^{g}\phi_{ij}^{g}).$ (48)
@@
$\bm{A}=\begin{bmatrix}\bm{A}_{\text{int}}\\ \bm{A}_{\text{bdy}}\end{bmatrix}\in\mathbb{R}^{N\times Q},\;\bm{b}=\begin{bmatrix}\bm{b}_{\text{int}}\\ \bm{b}_{\text{bdy}}\end{bmatrix}\in\mathbb{R}^{N},\;N=2N_{\text{int}}+N_{\text{bdy}},$ (55)
@@
${}^{11}\bm{A}_{\text{int}}^{k}=\begin{bmatrix}{}^{11}s^{\text{int},k}_{11}&\cdots&{}^{11}s^{\text{int},k}_{1J_{1}^{\rho}}&{}^{11}s^{\text{int},k}_{21}&\cdots&{}^{11}s^{\text{int},k}_{M^{\rho}J_{M^{\rho}}}\end{bmatrix},$ (58)
${}^{12}\bm{A}_{\text{int}}^{k}=\begin{bmatrix}{}^{12}s^{\text{int},k}_{11}&\cdots&{}^{12}s^{\text{int},k}_{1J_{1}^{g}}&{}^{12}s^{\text{int},k}_{21}&\cdots&{}^{12}s^{\text{int},k}_{M^{g}J_{M^{g}}}\end{bmatrix},$ (59)
${}^{21}\bm{A}_{\text{int}}^{k}=\begin{bmatrix}{}^{21}s^{\text{int},k}_{11}&\cdots&{}^{21}s^{\text{int},k}_{1J_{1}^{\rho}}&{}^{21}s^{\text{int},k}_{21}&\cdots&{}^{21}s^{\text{int},k}_{M^{\rho}J_{M^{\rho}}}\end{bmatrix},$ (60)
${}^{22}\bm{A}_{\text{int}}^{k}=\begin{bmatrix}{}^{22}s^{\text{int},k}_{11}&\cdots&{}^{22}s^{\text{int},k}_{1J_{1}^{g}}&{}^{22}s^{\text{int},k}_{21}&\cdots&{}^{22}s^{\text{int},k}_{M^{g}J_{M^{g}}}\end{bmatrix},$ (61)
${}^{31}\bm{A}_{\text{bdy}}^{k}=\begin{bmatrix}{}^{31}s^{\text{bdy},k}_{11}&\cdots&{}^{31}s^{\text{bdy},k}_{1J_{1}^{\rho}}&{}^{31}s^{\text{bdy},k}_{21}&\cdots&{}^{31}s^{\text{bdy},k}_{M^{\rho}J_{M^{\rho}}}\end{bmatrix},$ (62)
${}^{32}\bm{A}_{\text{bdy}}^{k}=\begin{bmatrix}{}^{32}s^{\text{bdy},k}_{11}&\cdots&{}^{32}s^{\text{bdy},k}_{1J_{1}^{g}}&{}^{32}s^{\text{bdy},k}_{21}&\cdots&{}^{32}s^{\text{bdy},k}_{M^{g}J_{M^{g}}}\end{bmatrix}.$ (63)
@@
$\bm{\rho}=\begin{bmatrix}\rho_{11}&\cdots&\rho_{1J_{1}^{\rho}}&\rho_{21}&\cdots&\rho_{M^{\rho}J_{M^{\rho}}}\end{bmatrix},$ (64)
@@
$\bm{g}=\begin{bmatrix}g_{11}&\cdots&g_{1J_{1}^{g}}&g_{21}&\cdots&g_{M^{g}J_{M^{g}}}\end{bmatrix},$ (65)
@@
Input: The set of collocation points $\{(\bm{x}_{\text{int}}^{k},\bm{v}_{\text{int}}^{k})\}_{1\leq k\leq N_{\text{int}}}$, $\{(\bm{x}_{\text{bdy}}^{k},\bm{v}_{\text{bdy}}^{k})\}_{1\leq k\leq N_{\text{bdy}}}$, $\{f_{\text{bdy}}^{k}\}_{1\leq k\leq N_{\text{bdy}}}$; The number of PoU functions $M^{\rho}$, $M^{g}$ and random feature functions $\{J_{i}^{\rho}\}_{1\leq i\leq M^{\rho}}$, $\{J_{i}^{g}\}_{1\leq i\leq M^{g}}$; The range of uniform distribution $[-B,B]$ for the initialization of neural network parameters; The Gauss-Legendre quadrature points and weights $\{(\omega_{j},\bm{v}_{j})\}_{j=1}^{N_{q}}$.

### v2

@@ 2.2 Random Feature Method
and $a_{{i+1},j}=b_{ij}$, which is analogous to the finite element method and one can generalize the partition to a more general shape if necessary. Assume $\bm{y}=(y_{1},\cdots,y_{d})^{T}\in\Omega$, we apply the following transformation to obtain a partition dependent normalized vector $\tilde{\bm{y}}_{i}=(\tilde{y}_{i1},\cdots,\tilde{y}_{id})^{T}$:
@@
where $\bm{w}_{ij}\in\mathbb{R}^{d}$ and $b_{ij}\in\mathbb{R}$ are randomly generated with the uniform distribution in $[-B,B]$ and fixed parameters, and $\sigma$ is a scalar function, named activation function. For simplicity, we assume that all $J_{i}$ are identical and denote them collectively as $J$. Generally, random feature functions are globally defined, whereas PDE solutions often exhibit local variations. To address this, RFM constructs multiple local solutions, each associated with a random feature model, and seamlessly integrates them using a partition of unity (PoU). The introduction of PoU generates local random feature functions, offering a more flexible and general strategy than domain decomposition or mesh generation. The number of partitions $M$ can be regarded as a mechanism for adapting to the solution’s local variations. The parameter $B$ is used to control the initialization range of the weights $\{\bm{w}_{ij}\}$ and $\{b_{ij}\}$.
@@ 2.3 Difficulty of RFM to resolve small scales
$\mathcal{T}^{\varepsilon}f_{M}=\mathcal{T}^{\varepsilon}\left(\sum_{i=1}^{M}\psi_{i}(x,v)\sum_{j=1}^{{J}}f_{ij}\phi_{ij}(x,v)\right)=\sum_{i=1}^{M}\sum_{j=1}^{{J}}f_{ij}\mathcal{T}^{\varepsilon}(\psi_{i}\phi_{ij}).$ (23)
@@
$\bm{A}=\begin{bmatrix}\bm{A}_{\text{int}}\\ \bm{A}_{\text{bdy}}\end{bmatrix}\in\mathbb{R}^{N\times{Z}},\;\bm{b}=\begin{bmatrix}\bm{b}_{\text{int}}\\ \bm{b}_{\text{bdy}}\end{bmatrix}\in\mathbb{R}^{N},\;N=N_{\text{int}}+N_{\text{bdy}},$ (24)
@@
$\bm{A}_{\text{int}}=\begin{bmatrix}s^{\text{int},1}_{11}&\cdots&s^{\text{int},1}_{1{J}}&s^{1}_{21}&\cdots&s^{\text{int},1}_{M{J}}\\ s^{\text{int},2}_{11}&\cdots&s^{\text{int},2}_{1{J}}&s^{2}_{21}&\cdots&s^{\text{int},2}_{M{J}}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ s^{\text{int},N_{\text{int}}}_{11}&\cdots&s^{\text{int},N_{\text{int}}}_{1{J}}&s^{\text{int},N_{\text{int}}}_{21}&\cdots&s^{\text{int},N_{\text{int}}}_{M{J}}\end{bmatrix},\,\bm{b}_{\text{int}}=-\varepsilon\begin{bmatrix}v_{\text{int}}^{1}\\ v_{\text{int}}^{2}\\ \vdots\\ v_{\text{int}}^{N_{\text{int}}}\end{bmatrix},$ (25)
$\bm{A}_{\text{bdy}}=\begin{bmatrix}s^{\text{bdy},1}_{11}&\cdots&s^{\text{bdy},1}_{1{J}}&s^{\text{bdy},1}_{21}&\cdots&s^{\text{bdy},1}_{M{J}}\\ s^{\text{bdy},2}_{11}&\cdots&s^{\text{bdy},2}_{1{J}}&s^{\text{bdy},2}_{21}&\cdots&s^{\text{bdy},2}_{M{J}}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ s^{\text{bdy},N_{\text{bdy}}}_{11}&\cdots&s^{\text{bdy},N_{\text{bdy}}}_{1{J}}&s^{\text{bdy},N_{\text{bdy}}}_{21}&\cdots&s^{\text{bdy},N_{\text{bdy}}}_{M{J}}\end{bmatrix},\;\bm{b}_{\text{bdy}}=\begin{bmatrix}f_{\text{bdy}}^{1}\\ f_{\text{bdy}}^{2}\\ \vdots\\ f_{\text{bdy}}^{N_{\text{bdy}}}\end{bmatrix}.$ (26)
@@
$\bm{f}=\begin{bmatrix}f_{11}&\cdots&f_{1{J}}&f_{21}&\cdots&f_{M{J}}\end{bmatrix}^{T},$ (27)
@@
First, we consider the problem (17) with $\varepsilon=1$ to validate the effectiveness of the RFM. Here, we set the number of partitions $M=1$ with the PoU function $\varphi_{b}$. We choose the $\tanh$ function as our activation function and the number of random feature functions $J=2^{n},n=3,4,5,6,7$. The weights $\bm{w}_{ij}$ and biases $b_{ij}$ are randomly generated with $B=1$. Besides, the collocation points are uniform grids with $(N_{x},N_{v})=(32,64)$. The integration of the operator $\mathcal{L}$ is approximated by Gauss-Legendre quadrature with $16$ points.
The relative $\ell^{2}$ error of the solution $f_{M}$ (as shown in (74)) concerning the degrees of freedom $Z$ is shown in Figure 2. One can observe that the error decreases exponentially to the degrees of freedom $Z$. We also plot the solution $f_{M}$ obtained by the random feature method with $J=2^{7}$ and reference solution $f_{\text{ex}}$ in Figure 3. The relative $\ell^{2}$ error of the RFM is $3.69\times 10^{-10}$.
Figure 2: Relative $\ell^{2}$ error of the RFM solution $f_{M}$ with respect to the degrees of freedom.
@@
Next, we test a sequential of problems (17) with $\varepsilon=10^{-2^{i}}$, $i=1,2,3,4$ with the same setting as the previous problem except for the number of random features ${J}$ and the collocation points $(N_{x},N_{v})$. We record the relative $\ell^{2}$ error of the RFM solution to the degrees of freedom and the number of collocation points in Table 1 and Table 2, respectively. One can find that the small-scale problems ($\varepsilon\ll 1$) are much more difficult to resolve than the large-scale problems in terms of the degrees of freedom and the number of collocation points.
Table 1: The dependency of the RFM on degrees of freedom. The collocation points are uniform grids with $(N_{x},N_{v})=(64,128)$.
$\varepsilon$ $J$
@@
Table 2: The dependency of the RFM on the number of collocation points. The number of random feature functions $J=128$.
@@
To resolve the multi-scale information, a natural consideration is to increase the number of degrees of freedom and the number of collocation points. The above strategy will lead to a large matrix $\bm{A}$ with a high condition number and make the least squares problem (28) ill-conditioned. Moreover, as shown in Table 2, when $\varepsilon$ is relatively large, for instance, $\varepsilon=10^{-2}$, increasing the number of collocation points can substantially enhance the accuracy. However, when $\varepsilon$ is exceedingly small, such as $\varepsilon=10^{-16}$, increasing the number of collocation points has a negligible impact and fails to significantly improve accuracy. Beyond the aspects previously discussed, we also investigated the impact of partition of unity (PoU) on this problem. The domain partitioning is denoted by $(M_{x},\,M_{v})$. The relative $\ell^{2}$ error are recorded in Table 3. It can be observed that when $\varepsilon\ll 1$, the effectiveness of the PoU is also quite limited. In some cases, it may even result in a decrease in accuracy. As can be seen from equation (29), relying solely on the original radiative transport equation imposes extremely stringent requirements on the mesh size of the partition.
Table 3: The dependency of the RFM on PoU. The collocation points are uniform grids with $(N_{x},N_{v})=(64,128)$ and the number of random feature functions $J=128$.
$\varepsilon$ $(M_{x},\,M_{v})$
(1, 1)11(1,\,1)( 1 , 1 ) (2, 1)21(2,\,1)( 2 , 1 ) (1, 2)12(1,\,2)( 1 , 2 ) (4, 1)41(4,\,1)( 4 , 1 ) (1, 4)14(1,\,4)( 1 , 4 )
[…]
The above analysis implies that the vanilla RFM based on least squares formulation for the radiative transfer equation (1) has tremendous difficulty in resolving small scales in the radiative transfer equation. This is not only due to the ill-conditioning of the least squares problem, but also the lack of capturing the small-scale part of the solution $f(x,v)$. Such a similar phenomenon is also observed in the numerical experiments in [28, 27] when solving the multiscale kinetic equations with small parameters through deep neural networks.
@@ 3 Asymptotic-Preserving Random Feature Method
Assume $\mathcal{F^{\varepsilon}}$ is the multi-scale model that depends on the scale parameter $\varepsilon$ and $\mathcal{F}^{0}$ is the corresponding asymptotic limit model as $\varepsilon\to 0$. Define $\mathcal{R}(\mathcal{F^{\varepsilon}})$ as the least-squares formulation of the model $\mathcal{F^{\varepsilon}}$ when solved using the random feature method. If $\mathcal{R}(\mathcal{F^{\varepsilon}})$ converges to $\mathcal{R}(\mathcal{F}^{0})$ as $\varepsilon\to 0$, and this limit is precisely the least-squares formulation of the limit model $\mathcal{F}^{0}$, then the method is called asymptotic-preserving.
@@
Therefore, equations (34) and (36) constitute the micro-macro system of the radiative transfer equation (31):
@@
respectively. Here, $\{(\bm{x}_{\text{int}}^{k},\bm{v}_{\text{int}}^{k})\}_{1\leq k\leq N_{\text{int}}}$ and $\{(\bm{x}_{\text{bdy}}^{k},\bm{v}_{\text{bdy}}^{k})\}_{1\leq k\leq N_{\text{bdy}}}$ are the interior and boundary collocation points, and $\{f_{\text{bdy}}^{k}\}_{1\leq k\leq N_{\text{bdy}}}$ are the boundary values. $\{\lambda_{\text{int},1}^{k}\},\{\lambda_{\text{int},2}^{k}\}$ and $\{\lambda_{\text{bdy}}^{k}\}$ are regularization parameters. The setting of the regularization parameters will be determined through a simple but effective way in (69). The degrees of freedom of the APRFM are $Z=M^{\rho}J^{\rho}+M^{g}J^{g}$.
Taking $\varepsilon\to 0$, the least squares problem (41) reduces to
$\mathcal{R}_{\text{int}}^{0}$ $=\sum_{k=1}^{N_{\text{int}}}\lambda_{\text{int},1}^{k}|\left\langle\bm{v}\cdot\nabla_{\bm{x}}g_{M}\right\rangle(\bm{x}_{\text{int}}^{k})+\sigma_{a}(\bm{x}_{\text{int}}^{k})\rho_{M}(\bm{x}_{\text{int}}^{k})-\left\langle Q\right\rangle(\bm{x}_{\text{int}}^{k})|^{2}$ (43)
$+\sum_{k=1}^{N_{\text{int}}}\lambda_{\text{int},2}^{k}|\bm{v}_{\text{int}}^{k}\cdot\nabla_{\bm{x}}\rho_{M}(\bm{x}_{\text{int}}^{k})-\sigma_{s}(\bm{x}_{\text{int}}^{k})\mathcal{L}g_{M}(\bm{x}_{\text{int}}^{k},\bm{v}_{\text{int}}^{k})|^{2},$
which corresponds to the least squares formulation of the system
$\begin{cases}\left\langle\bm{v}\cdot\nabla_{\bm{x}}g\right\rangle+\sigma_{a}\rho=\left\langle Q\right\rangle,\\ \bm{v}\cdot\nabla_{\bm{x}}\rho=\sigma_{s}\mathcal{L}g.\end{cases}$ (44)
From the second equation of (44), we obtain $g=\mathcal{L}^{-1}(\frac{1}{\sigma_{s}(\bm{x})}\bm{v}\cdot\nabla_{\bm{x}}\rho)$. Substituting this into the first equation yields the limiting equation (5). Thus, this proposed method is asymptotic-preserving.
@@
$\mathcal{T}_{11}^{\varepsilon}\rho_{M}=\sum_{i=1}^{M^{\rho}}\sum_{j=1}^{{J^{\rho}}}\rho_{ij}\mathcal{T}_{11}^{\varepsilon}(\psi_{i}^{\rho}\phi_{ij}^{\rho}),\;\mathcal{T}_{12}^{\varepsilon}g_{M}=\sum_{i=1}^{M^{g}}\sum_{j=1}^{{J^{g}}}g_{ij}\mathcal{T}_{12}^{\varepsilon}(\psi_{i}^{g}\phi_{ij}^{g}),$ (49)
@@
$\mathcal{T}_{21}^{\varepsilon}\rho_{M}=\sum_{i=1}^{M^{\rho}}\sum_{j=1}^{{J^{\rho}}}\rho_{ij}\mathcal{T}_{21}^{\varepsilon}(\psi_{i}^{\rho}\phi_{ij}^{\rho}),\;\mathcal{T}_{22}^{\varepsilon}g_{M}=\sum_{i=1}^{M^{g}}\sum_{j=1}^{{J^{g}}}g_{ij}\mathcal{T}_{22}^{\varepsilon}(\psi_{i}^{g}\phi_{ij}^{g}).$ (50)
@@
$\bm{A}=\begin{bmatrix}\bm{A}_{\text{int}}\\ \bm{A}_{\text{bdy}}\end{bmatrix}\in\mathbb{R}^{N\times Z},\;\bm{b}=\begin{bmatrix}\bm{b}_{\text{int}}\\ \bm{b}_{\text{bdy}}\end{bmatrix}\in\mathbb{R}^{N},\;N=2N_{\text{int}}+N_{\text{bdy}},$ (57)
@@
${}^{11}\bm{A}_{\text{int}}^{k}=\begin{bmatrix}{}^{11}s^{\text{int},k}_{11}&\cdots&{}^{11}s^{\text{int},k}_{1{J^{\rho}}}&{}^{11}s^{\text{int},k}_{21}&\cdots&{}^{11}s^{\text{int},k}_{M^{\rho}{J^{\rho}}}\end{bmatrix},$ (60)
${}^{12}\bm{A}_{\text{int}}^{k}=\begin{bmatrix}{}^{12}s^{\text{int},k}_{11}&\cdots&{}^{12}s^{\text{int},k}_{1{J^{g}}}&{}^{12}s^{\text{int},k}_{21}&\cdots&{}^{12}s^{\text{int},k}_{M^{g}{J^{g}}}\end{bmatrix},$ (61)
${}^{21}\bm{A}_{\text{int}}^{k}=\begin{bmatrix}{}^{21}s^{\text{int},k}_{11}&\cdots&{}^{21}s^{\text{int},k}_{1{J^{\rho}}}&{}^{21}s^{\text{int},k}_{21}&\cdots&{}^{21}s^{\text{int},k}_{M^{\rho}{J^{\rho}}}\end{bmatrix},$ (62)
${}^{22}\bm{A}_{\text{int}}^{k}=\begin{bmatrix}{}^{22}s^{\text{int},k}_{11}&\cdots&{}^{22}s^{\text{int},k}_{1{J^{g}}}&{}^{22}s^{\text{int},k}_{21}&\cdots&{}^{22}s^{\text{int},k}_{M^{g}{J^{g}}}\end{bmatrix},$ (63)
${}^{31}\bm{A}_{\text{bdy}}^{k}=\begin{bmatrix}{}^{31}s^{\text{bdy},k}_{11}&\cdots&{}^{31}s^{\text{bdy},k}_{1{J^{\rho}}}&{}^{31}s^{\text{bdy},k}_{21}&\cdots&{}^{31}s^{\text{bdy},k}_{M^{\rho}{J^{\rho}}}\end{bmatrix},$ (64)
${}^{32}\bm{A}_{\text{bdy}}^{k}=\begin{bmatrix}{}^{32}s^{\text{bdy},k}_{11}&\cdots&{}^{32}s^{\text{bdy},k}_{1{J^{g}}}&{}^{32}s^{\text{bdy},k}_{21}&\cdots&{}^{32}s^{\text{bdy},k}_{M^{g}{J^{g}}}\end{bmatrix}.$ (65)
@@
$\bm{\rho}=\begin{bmatrix}\rho_{11}&\cdots&\rho_{1{J^{\rho}}}&\rho_{21}&\cdots&\rho_{M^{\rho}{J^{\rho}}}\end{bmatrix},$ (66)
@@
$\bm{g}=\begin{bmatrix}g_{11}&\cdots&g_{1{J^{g}}}&g_{21}&\cdots&g_{M^{g}{J^{g}}}\end{bmatrix},$ (67)
@@
To reduce the condition number of the least squares problem, we rescale the residuals of each term in (68) based on the maximum absolute value, ensuring that their magnitudes are on the same order. Specifically, the matrix $\bm{A}=\left(a_{kj}\right)\in\mathbb{R}^{N\times Z}$, and the vector $\bm{b}=[b_{1},\cdots,b_{N}]^{T}\in\mathbb{R}^{N}$ are redefined and the regularization parameters $\lambda^{k}\in\{\lambda_{\text{int},1}^{k},\lambda_{\text{int},2}^{k},\lambda_{\text{bdy}}^{k}\}$ can be defined as follows:
$\lambda^{k}=\frac{1}{\max_{1\leq j\leq Z}|a_{kj}|},\;k=1,\cdots,N.$ (69)
Correspondingly, the elements of the matrix $\bm{A}$ and the vector $\bm{b}$ in (40) should be transformed, with their respective elements given by
$a_{kj}=\lambda^{k}a_{kj},b_{k}=\lambda^{k}b_{k},\;k=1,\cdots,N,j=1,\cdots,Z.$ (70)
@@
Input: The set of collocation points $\{(\bm{x}_{\text{int}}^{k},\bm{v}_{\text{int}}^{k})\}_{1\leq k\leq N_{\text{int}}}$, $\{(\bm{x}_{\text{bdy}}^{k},\bm{v}_{\text{bdy}}^{k})\}_{1\leq k\leq N_{\text{bdy}}}$, $\{f_{\text{bdy}}^{k}\}_{1\leq k\leq N_{\text{bdy}}}$; The number of PoU functions $M^{\rho}$, $M^{g}$ and random feature functions $J^{\rho},J^{g}$; The range of uniform distribution $[-B,B]$ for the initialization of neural network parameters; The Gauss-Legendre quadrature points and weights $\{(\omega_{j},\bm{v}_{j})\}_{j=1}^{N_{q}}$.